# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

### v1

@@ IV PTM analysis for general quantum channels
    Decide which types of features $\phi_{r}$ are required to learn the dataset. For example, learning the flow map of a Markovian dynamical system typically requires low-order polynomial features. For harmonic time series and non-Markovian data, linear delay features are needed to capture temporal dependencies. Results like in Table 1 can aid systematic reservoir design.
@@
    Select suitable measurement observables $P_{k}$. Ideally, the Heisenberg-evolved observables $P_{k}^{\text{(H)}}(t)$ have strong overlap with the observable $P_{r}$ generating the desired feature $\phi_{r}$ [see Eq. 17]. One possibility to achieve this is to determine an optimal measurement operator $M^{*}$ via primal or dual optimization, which can then be further expanded into a set of rotated Pauli-Z operators [29]. Another possibility is temporal multiplexing [1, 44, 60, 61], discussed further in the PTM framework below (see Section IV.2).
@@
In temporal multiplexing [62, 1], instead of working at a fixed evolution time $t$ and a potentially large set of observables $P_{k}$, one selects a small subset $\mathcal{S}$ of size ${m}=|\mathcal{S}|<d^{2}$ and evaluate them at several times $t_{1},\ldots,t_{L}$. Correspondingly, we stack a subset $\mathcal{S}$ of rows of $V(t)$ into a ${m}{L}\times d^{2}$ observability matrix
$R=\begin{pmatrix}\mathcal{S}V(t_{1})\\ \mathcal{S}V(t_{2})\\ \vdots\\ \mathcal{S}V(t_{L}).\end{pmatrix}$ (24)
The case without temporal multiplexing corresponds to L=1{L}=1. For $\mathcal{S}=\mathbbm{1}_{d^{2}}$ and L=1{L}=1, one then recovers the full unital PTM $R=V(t)$.
@@ IV.1 Classical representation
According to Eq. 16, a QELM has an equivalent classical representation as a PTM-weighted nonlinear function library,
$f(\mathbf{u})=\mathbf{w}^{\top}R\boldsymbol{\upphi}(\mathbf{u})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{u})$ (25)
where $R\in\mathbb{R}^{{B}\times{q}}$ is the effective PTM (24) (including any row-selection and temporal multiplexing) and $F_{k}=R_{kj}\phi_{j}$ are the Pauli features available at the readout. We denote the effective number of features by ${q}\leq d^{2}$, excluding features identically vanishing due to the encoding, while B{B} is the total number of measurements, e.g. B=mL{B}={m}{L}.
@@
$\mathbf{w}^{*}=(GG^{\top}+\lambda\mathbbm{1}_{B})^{-1}G\mathbf{y}=R(CR^{\top}R+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y},$ (26)
@@
$G\equiv[\mathbf{F}(\mathbf{u}_{1})\cdots\mathbf{F}(\mathbf{u}_{P})]=R\Phi\in\mathbb{R}^{{B}\times P},\quad GG^{\top}=RCR^{\top}\in\mathbb{R}^{{B}\times{B}}$ (28)
@@
$\mathbf{w}^{*}=V(C+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y}.$ (29)
Moreover, since $V$ is orthogonal, it drops out from the predictor (25), consistent with the absence of a reservoir unitary in the kernel-representation (63). Such a QELM is then simply a regressor acting on the full $d^{2}$-dimensional feature library $\boldsymbol{\upphi}$. In the context of time series processing (especially when using some inputs $u_{j}$ to encode delay states) this is also known as a “next-generation” reservoir computer (NG-RC) [64] or, equivalently, a nonlinear vector autoregressive process (NVAR). When $R$ merely has full column rank, it can be interpreted to induce a regressor operating on the mixed feature set $\boldsymbol{\upphi}^{\prime}=R\boldsymbol{\upphi}$.
@@
$\mathrm{rank}(G)=\min(\mathrm{rank}(R),\mathrm{rank}(\Phi))\leq\min({B},P,{q}).$ (30)
For sufficiently diverse data and large sample number $P$, $\Phi$ typically achieves a rank near its maximum $\min({q},P)$. However, it can suffer from severe ill-conditioning since the features $\phi_{k}$ are in general not orthogonal. Anticipating results discussed below (see Fig. 4), for random Hamiltonians $\mathrm{rank}(R)$ typically grows monotonically with the measurement budget B{B} and saturates at $\mathrm{rank}(\Phi)\sim{q}$. By contrast, for structured Hamiltonians like the TFIM, $\mathrm{rank}(R)$ and thus $\mathrm{rank}(G)$ can saturate at a value significantly smaller than q{q}.
@@ IV.2 Krylov growth and Pauli weight spreading
Temporal multiplexing effectively “reverses” the information scrambling caused by quantum evolution. The latter is captured by the Krylov growth of an operator under conjugation [65, 59, 60], which we summarize here in our setting. Specifically, the unitary Heisenberg-evolution [Eq. 18] of a given Pauli observable $P_{k}$ satisfies
$P_{k}^{\text{(H)}}(t)\equiv U(t)^{\dagger}P_{k}U(t)=e^{iHt}P_{k}e^{-iHt}=\sum_{\ell=0}^{\infty}\frac{(it)^{\ell}}{\ell!}\,\mathrm{ad}_{H}^{\ell}(P_{k})$ (31)
@@
$(Z_{j_{1}}\cdots Z_{j_{k}})^{\text{H}}(t)\big|_{\text{zz-x}}=U_{\text{H}}^{\dagger}(X_{j_{1}}\cdots X_{j_{k}})^{\text{H}}(t)\big|_{\text{xx-z}}U_{\text{H}}$ (33)
@@
$P_{k}^{\text{(H)}}(t)=\sum_{j=0}^{d^{2}-1}V_{kj}(t)P_{j}$ (34)
where we used cyclicity of the trace to introduce $V(t)$ from Eq. 19. Thus, for non-Clifford $H$, an initial weight fully localized on $P_{k}$ spreads to other (higher and lower order) Pauli strings as $t$ increases. A convenient diagnostic is the Pauli (Hamming) weight $\nu(P_{j})\in\{0,\ldots,n\}$, i.e., the number of non-identity factors in $P_{j}$, and its coefficient-weighted average
@@
For local (including nonintegrable/chaotic ones) Hamiltonians, the Lieb–Robinson bound [67, 68, 69, 70, 47] implies that an initially local operator can only develop support within distance $\lesssim v_{\mathrm{LR}}t$ until finite-size saturation:
@@
with a typical late-time value $\bar{\nu}_{k}\approx\tfrac{3}{4}n$, corresponding to a uniform distribution of coefficients $V_{kj}^{2}$ over Pauli strings. From the definition (18) and the expansion (34), it straightforwardly follows that $\|P_{k}^{(H)}\|=d=d\sum_{j=0}^{d^{2}-1}\|V_{kj}\|^{2}$, and thus $\sum_{j=0}^{d^{2}-1}V_{kj}^{2}=1$ for all $t$. Consequently, for Haar-random unitaries, one expects a typical late-time magnitude
$|V_{kj}(t\to\infty)|=2^{-n}$ (37)
This is the origin of the exponential concentration effect, where the output of the QELM essentially looses any information about the input for large qubit numbers $n\to\infty$ [51]. Integrable models (such as the TFIM at its integrable points) typically show more structured behavior due to stable quasiparticle modes. An extreme case is a Clifford unitary, for which $V_{kj}=\mathrm{const}.$ By contrast, a Haar-random unitary on the full Hilbert space can generate rapid delocalization [71, 72, 73, 74, 75]. The evolution of the Pauli weight average is illustrated in Fig. 1.
@@ IV.3 Pauli feature decoding
Following Eq. 25, we denote the set of Pauli features available at the readout by $F_{k}=R_{kj}\phi_{j}$ ($k=1,\ldots,{m}$), which emerge by linear transformations of the input features $\phi_{j}$ ($j=1,\ldots,{q}\leq d^{2}$) via the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$ [Eq. 24], possibly including temporal multiplexing. We now ask for conditions that enable the regressor to isolate a given input feature $\phi_{r}(x)$ by finding weights $\mathbf{w}$ such that (approximately)
$\mathbf{w}\cdot\mathbf{F}\propto\phi_{r}\quad\Leftrightarrow\quad\mathbf{w}^{\top}R=\omega\mathbf{e}_{r}^{\top}$ (38)
@@
If $R$ corresponds to the full unital channel, then $R=V(t)\in\mathbb{R}^{{q}\times{q}}$ is invertible and any feature can be decoded by
$\mathbf{w}=\omega V(t)\mathbf{e}_{r}.\qquad\text{(unital channel)}$ (39)
@@
$\gamma_{r}^{2}\equiv(R^{+}R)_{rr},\qquad\sum_{r=1}^{d^{2}}\gamma_{r}^{2}=\mathrm{rank}(R)\leq d^{2}$ (40)
@@
$\mathbf{w}=\omega\left((R^{+})_{r,:}\right)^{\top}.$ (41)
These coincide with the minimum-norm least-squares solution. Exact decodability of the $r$-th feature is indicated by $\gamma_{r}^{2}=1$. A values $\gamma_{r}^{2}<1$ implies that the feature can only be reconstructed up to an error $1-\gamma_{r}^{2}$. We take $\gamma_{r}^{2}\gtrsim 0.5$ as a useful decodability threshold. Note that $\gamma_{r}^{2}$ is a geometric notion of feature decodability and ignores statistical correlations (see Appendix C for further discussion).
Both score and weights preserve tensor product structure of the general PTM $R$: if $R=\otimes_{i=1}^{n}R_{i}$ then $\gamma_{r}^{2}=\prod_{i=1}^{n}\gamma_{r,i}^{2}$ and $\mathbf{w}=\otimes_{i=1}^{n}\mathbf{w}_{i}$. For $R=\mathcal{S}V(t)$, one obtains $\gamma_{r}^{2}=\sum_{k\in\mathcal{S}}V(t)^{2}_{kr}$ showing that feature $r$ is decodable only to the extent that the “spread” of that Pauli under $V(t)$ lands in the measured subset $\mathcal{S}$. [^2] If $R$ contains a $t=0$ block, then since $V(0)=\mathbbm{1}_{d^{2}}$ the corresponding rows of $R$ include $\mathbf{e}_{j}^{\top}$ for all $j\in\mathcal{S}$ (up to basis ordering), hence $\mathbf{e}_{j}\in\mathrm{row}(R)$. Since $R^{+}R$ is the orthogonal projector onto $\mathrm{row}(R)$, $(R^{+}R)\mathbf{e}_{j}=\mathbf{e}_{j}$ and thus $\gamma_{j}^{2}=(R^{+}R)_{jj}=1$. Specifically, for Haar-random unitaries, late-time decodability decays exponentially with qubit number $n$ [see Eq. 37],
$\gamma_{r}^{2}\big|_{t\to\infty}\sim|\mathcal{S}|4^{-n}$ (42)
Noisy channels generically have invertible PTMs [see Section D.2] and hence do not strictly reduce the accessible feature subspace. However, they are contracting and can effectively suppress high-weight Pauli features [see Eq. 84], requiring large weight norms $\|\mathbf{w}\|$ and thereby degrading conditioning. Such non-CPTP maps are generally unstable under finite sampling and thus effectively lead to information loss.
@@
The evolution with multiplexing steps L{L} of the mean decodability $\bar{\gamma}_{r}^{2}$ of each Pauli sector is further illustrated in Fig. 4. Since we now start with $t=1$, the selected low-weight Paulis are already spread and initial decodability is low on average  [^3] Including $t\approx 0$ in the temporal multiplexing scheme improves decodability at early iterations, but does not affect the final decodability for sufficiently large L{L}.. Compared to structured Hamiltonians, like the TFIM, random Hamiltonians lead to a consistently higher rank of $R$ and a more even decodability over Pauli sectors for large L{L}. In general, we find that temporal multiplexing with Pauli $Z$ and $ZZ$ observables and random unitaries requires ${L}\sim d^{2}/|\mathcal{S}|$ measurements to achieve high decoding scores in all Pauli sectors, along with a saturated rank $\mathrm{rank}(R)\sim|\mathcal{S}|{L}\sim d^{2}$. In practice, however, this exponential measurement budget is usually not needed, since important features can be statistically decoded with much fewer measurements (see Figs. 6 and 7 below).
@@
The decodability framework is essentially a reformulation of pseudo-inverse solution of the RC training and provides a handle on the available features [78, 79, 80, 52].The decodability score (40) is a data-independent property of the chosen measurement observables and quantum operations after the encoding. For moderate qubit numbers $n$, full Pauli feature decodability is guaranteed when one either performs a full measurement (${m}=d^{2}$) on an invertible quantum channel, or one employs a temporal multiplexing scheme leading to a full column rank of $R$, again requiring an exponential measurement budget ${m}{L}\sim d^{2}$. Practically sufficient decodability can already be achieved with modest scores and polynomial-in-$n$ measurement budgets, as typically done in applications [1, 44, 60, 61]. Structured unitaries can be advantageous in the large-$n$ limit to avoid exponential concentration issues [51].

### v2

@@ IV PTM analysis for a unitary quantum channel
    Decide which types of features $\{\phi_{r}\}$ are best suited to learn the dataset. For example, non-Markovian time series (including harmonic signals) require delay coordinates [65, 66], whereas Markovian dynamical systems can already be modeled by instantaneous monomial features. While commonly used encodings support universal function approximation [26, 6, 27, 28, 29, 30], they can differ in their approximation errors, generalization behavior, or robustness during autoregressive rollout.
@@
    Select suitable measurement observables $\{P_{k}\}$. Ideally, the Heisenberg-evolved observables $P_{k}^{\text{(H)}}(t)$ have strong overlap with the observable $P_{r}$ generating the desired feature $\phi_{r}$ [see Eq. 17]. One possibility to achieve this is to determine an optimal measurement operator $M^{*}$ via primal or dual optimization, which can then be further expanded into a set of rotated Pauli-Z operators [32]. Another possibility is temporal multiplexing [1, 47, 67, 68], discussed further in the PTM framework below (see Section IV.2).
@@
In temporal multiplexing [69, 1], instead of working at a fixed evolution time $t$ and a potentially large set of observables $P_{k}$, one selects a small subset $\mathcal{S}$ of size m{m} and evaluates them at several times $t_{1},\ldots,t_{L}$ [cf. Eq. 16]. One then stacks the corresponding rows of $V(t)$ into a ${m}{L}\times d^{2}$ observability matrix
$R=\begin{pmatrix}SV(t_{1})\\ SV(t_{2})\\ \vdots\\ SV(t_{L})\end{pmatrix}$. (24)
The case without temporal multiplexing corresponds to L=1{L}=1. The full unitary PTM $R=V(t)$ is then recovered for $S=\mathbbm{1}_{d^{2}}$.
@@ IV.1 Classical representation
Let ${q}\leq d^{2}$ be the effective number of features (excluding features identically vanishing due to the encoding), and B{B} be the total number of measurements, e.g. B=mL{B}={m}{L}. According to Eq. 16, a QELM has an equivalent classical representation as a PTM-weighted nonlinear function library,
$f(\mathbf{u})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{u}),\qquad\mathbf{F}(\mathbf{u})\equiv R\boldsymbol{\upphi}(\mathbf{u}),$ (25)
where $R\in\mathbb{R}^{{B}\times{q}}$ is the effective PTM (24) (including any row-selection and temporal multiplexing) and $F_{k}=R_{kj}\phi_{j}$ are the Pauli features available at the readout.
@@
$\hat{\mathbf{w}}=(GG^{\top}+\lambda\mathbbm{1}_{B})^{-1}G\mathbf{y}=R(CR^{\top}R+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y},$ (26)
@@
$G\equiv[\mathbf{F}(\mathbf{u}_{1})\cdots\mathbf{F}(\mathbf{u}_{P})]=R\Phi\in\mathbb{R}^{{B}\times P},\quad GG^{\top}=RCR^{\top}\in\mathbb{R}^{{B}\times{B}}.$ (28)
@@
$\hat{\mathbf{w}}=V(C+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y}.$ (29)
Moreover, since $V$ is orthogonal, it drops out from the predictor (25), consistent with the absence of a reservoir unitary in the kernel-representation (59). Such a QELM is then simply a regressor acting on the full $d^{2}$-dimensional feature library $\boldsymbol{\upphi}$. In the context of time series processing (especially when using some inputs $u_{j}$ to encode delay states) this is reduces to a “next-generation” reservoir computer (NG-RC) [71] or, equivalently, a nonlinear vector autoregressive process (NVAR). When $R$ merely has full column rank, it can be interpreted to induce a regressor operating on the mixed feature set $\boldsymbol{\upphi}^{\prime}=R\boldsymbol{\upphi}$.
@@
$\mathrm{rank}(G)\leq\min(\mathrm{rank}(R),\mathrm{rank}(\Phi))\leq\min({B},P,{q}).$ (30)
For sufficiently diverse data, $\Phi$ typically achieves a rank near its maximum $\min({q},P)$, but can suffer from ill-conditioning since the features $\phi_{k}$ are in general not orthogonal. Accordingly, for large sample number $P$, the rank of the readout approximately follows the rank of the effective PTM:
$\mathrm{rank}(G)\simeq\mathrm{rank}(R).$ (31)
For random Hamiltonians, $\mathrm{rank}(R)$ typically grows proportionally to the measurement budget B{B} and then saturates near q{q} (see Fig. 4). By contrast, structured Hamiltonians like the TFIM are not fully scrambling, resulting in $\mathrm{rank}(R)\ll{q}$. The achieved ranks indicate the quality of the temporal multiplexing scheme and thus of the training accuracy of the QELM (see Fig. 10).
@@ IV.2 Krylov growth and Pauli weight spreading
Temporal multiplexing enables the regressor to “reverse” the information scrambling caused by quantum evolution. The latter is captured by the Krylov growth of an operator under conjugation [72, 64, 67], which we summarize here in our setting. Specifically, the unitary Heisenberg-evolution [Eq. 18] of a given Pauli observable $P_{k}$ satisfies
$P_{k}^{\text{(H)}}(t)\equiv U(t)^{\dagger}P_{k}U(t)=e^{iHt}P_{k}e^{-iHt}=\sum_{\ell=0}^{\infty}\frac{(it)^{\ell}}{\ell!}\,\mathrm{ad}_{H}^{\ell}(P_{k}),$ (32)
@@
$(Z_{j_{1}}\cdots Z_{j_{k}})^{\text{H}}(t)\big|_{\text{zz-x}}=U_{\text{H}}^{\dagger}(X_{j_{1}}\cdots X_{j_{k}})^{\text{H}}(t)\big|_{\text{xx-z}}U_{\text{H}},$ (34)
@@
$P_{k}^{\text{(H)}}(t)=\sum_{j=0}^{d^{2}-1}V_{kj}(t)P_{j},$ (35)
where we used cyclicity of the trace to introduce $V(t)$ from Eq. 19. Thus, for non-Clifford $H$, an initial weight fully localized on $P_{k}$ can spread to other (higher and lower order) Pauli strings as $t$ increases. A convenient diagnostic is the Pauli (Hamming) weight $\nu(P_{j})\in\{0,\ldots,n\}$, i.e., the number of non-identity factors in $P_{j}$, and its coefficient-weighted average
@@
For local (including nonintegrable/chaotic ones) Hamiltonians, the Lieb–Robinson bound [74, 75, 76, 77, 50] heuristically implies that an initially local operator can only develop support within distance $\lesssim v_{\mathrm{LR}}t$ until finite-size saturation:
@@
with a typical late-time value $\bar{\nu}_{k}\approx\tfrac{3}{4}n$, corresponding to a uniform distribution of coefficients $V_{kj}^{2}$ over Pauli strings. From the definition (18) and the expansion (35), it straightforwardly follows that $\|P_{k}^{(H)}\|^{2}=d=d\sum_{j=0}^{d^{2}-1}V_{kj}^{2}$, and thus $\sum_{j=0}^{d^{2}-1}V_{kj}^{2}=1$ for all $t$. Consequently, for Haar-random unitaries, one expects a typical late-time magnitude
$|V_{kj}(t\to\infty)|=2^{-n}.$ (38)
This is the origin of the exponential concentration effect, which makes accessing information about the input infeasible for large qubit numbers $n\to\infty$ [53]. Integrable models (such as the TFIM at its integrable points) typically show more structured behavior due to stable quasiparticle modes. An extreme case is a Clifford unitary, for which $V_{kj}$ is a signed permutation matrix (no spreading). By contrast, a Haar-random unitary on the full Hilbert space can generate rapid delocalization [78, 79, 80, 81, 82]. The evolution of the Pauli weight average is illustrated in Fig. 1.
@@ IV.3 Pauli feature decoding
Following Eq. 25, we denote the set of Pauli features available at the readout by $F_{k}=R_{kj}\phi_{j}$ ($k=1,\ldots,{B}$), which emerge by linear transformations of the input features $\phi_{j}$ ($j=1,\ldots,{q}\leq d^{2}$) via the effective PTM $R\in\mathbb{R}^{{B}\times{q}}$ [Eq. 24], possibly including temporal multiplexing. We now ask for conditions that enable the regressor to isolate a given input feature $\phi_{r}(\mathbf{x})$ by finding weights $\mathbf{w}$ such that (approximately)
$\mathbf{w}\cdot\mathbf{F}\propto\phi_{r}\quad\Leftrightarrow\quad\mathbf{w}^{\top}R=\omega\mathbf{e}_{r}^{\top},$ (39)
@@
If $R$ corresponds to the full unitary channel, then $R=V(t)\in\mathbb{R}^{{q}\times{q}}$ is invertible and any feature can be decoded by
$\mathbf{w}=\omega V(t)\mathbf{e}_{r}.\qquad\text{(unitary channel)}$ (40)
@@
$\gamma_{r}^{2}\equiv(R^{+}R)_{rr},\qquad\sum_{r=1}^{{q}}\gamma_{r}^{2}=\mathrm{rank}(R)\leq{q},$ (41)
@@
$\mathbf{w}=\omega\left((R^{+})_{r,:}\right)^{\top},$ (42)
in terms of the Moore–Penrose pseudoinverse $R^{+}$. These coincide with the minimum-norm least-squares solution. Exact decodability of the $r$-th feature is indicated by $\gamma_{r}^{2}=1$. A value $\gamma_{r}^{2}<1$ implies that the feature can only be reconstructed up to an error $\sim\sqrt{1-\gamma_{r}^{2}}\|\phi\|_{2}$. We take $\gamma_{r}^{2}\gtrsim 0.5$ as a useful decodability threshold. Note that $\gamma_{r}^{2}$ is a geometric notion and ignores statistical correlations or noise floor (see Appendix C for further discussion).
Both score and weights preserve tensor product structure of the general PTM $R$: if $R=\otimes_{i=1}^{n}R_{i}$ then $\gamma_{r}^{2}=\prod_{i=1}^{n}\gamma_{r,i}^{2}$ and $\mathbf{w}=\otimes_{i=1}^{n}\mathbf{w}_{i}$. For $R=SV(t)$, one obtains $\gamma_{r}^{2}=\sum_{k\in\mathcal{S}}V(t)^{2}_{kr}$, showing that feature $r$ is decodable only to the extent that the “spread” of that Pauli under $V(t)$ lands in the measured subset $\mathcal{S}$. [^2] If $R$ contains a $t=0$ block, then since $V(0)=\mathbbm{1}_{d^{2}}$ the corresponding rows of $R$ include $\mathbf{e}_{j}^{\top}$ for all $j\in\mathcal{S}$ (up to basis ordering), hence $\mathbf{e}_{j}\in\mathrm{row}(R)$. Since $R^{+}R$ is the orthogonal projector onto $\mathrm{row}(R)$, $(R^{+}R)\mathbf{e}_{j}=\mathbf{e}_{j}$ and thus $\gamma_{j}^{2}=(R^{+}R)_{jj}=1$. Specifically, for Haar-random unitaries, late-time decodability decays exponentially with qubit number $n$ [see Eq. 38],
$\gamma_{r}^{2}\big|_{t\to\infty}\sim|\mathcal{S}|4^{-n}.$ (43)
Noisy channels can have invertible PTMs [see Section D.2] and then do not strictly reduce the accessible feature subspace. If they are contracting, they effectively suppress high-weight Pauli features [see Eq. 81], but require large weight norms $\|\mathbf{w}\|$ which degrades conditioning. Such maps are unstable under finite sampling and thus effectively lead to information loss.
@@
The evolution with multiplexing steps L{L} of the mean decodability $\bar{\gamma}_{r}^{2}$ of each Pauli sector is further illustrated in Fig. 4. Since we now start with $t=1$, the selected low-weight Paulis are already spread and initial decodability is low on average  [^3] Including $t\approx 0$ in the temporal multiplexing scheme improves decodability at early iterations, but does not affect the final decodability for sufficiently large L{L}.. Structured Hamiltonians, like the TFIM [Eq. 5], can have strong impact on the achievable decodability score depending on the variant or the specific observables used. Comparing Fig. 4(c) and (g), Pauli-$Z$’s evolving under the Hamiltonian $H_{\text{TFIM}}^{\text{xx-z}}$ render only weak scores compared to the Hamiltonian $H_{\text{TFIM}}^{\text{zz-x}}$  [^4] By global Hadamard equivalence, the statement flips for Pauli-$X$ as initial operators.. A more detailed analysis using the Jordan-Wigner/Majorana formalism [86, 87] shows that, under $H_{\text{TFIM}}^{\text{xx-z}}$, the Pauli strings generated from an initial $Z$ are mainly consisting of $Z$-strings with XX/$Y$ endpoints, of which there are $\sim\mathcal{O}(n^{2})$. By contrast, spreading under $H_{\text{TFIM}}^{\text{zz-x}}$ almost fully saturates the Pauli basis ($\mathcal{O}(4^{n})$ strings, up to sub-exponential factors).
Random Hamiltonians, by contrast, typically populate the full Pauli basis. Under temporal multiplexing they thus consistently lead to a higher rank of $R$ and a more even decodability over Pauli sectors for large L{L}. In general, we find that temporal multiplexing with Pauli $Z$ and $ZZ$ observables and random unitaries requires ${L}\sim d^{2}/|\mathcal{S}|$ measurements to achieve high decoding scores in all Pauli sectors, along with a saturated rank $\mathrm{rank}(R)\sim|\mathcal{S}|{L}\sim d^{2}$. In practice, however, this exponential measurement budget is rarely needed, since relevant features can be statistically decoded with much fewer measurements (see Figs. 6 and 7 below).
@@
The decodability framework reinterprets the pseudo-inverse solution of the RC training in terms of individually decodable features [88, 89, 90, 54]. The decodability score (41) is a data-independent property of the chosen measurement observables and quantum operations after the encoding. For moderate qubit numbers $n$, full Pauli feature decodability is guaranteed when one either performs a full measurement (${m}=d^{2}$) on an invertible quantum channel, or one employs a temporal multiplexing scheme leading to a full column rank of $R$, again requiring an exponential measurement budget ${m}{L}\sim d^{2}$. Practically sufficient decodability can already be achieved with modest scores and polynomial-in-$n$ measurement budgets, as typically done in applications [1, 47, 67, 68]. Structured unitaries can be advantageous in the large-$n$ limit to avoid exponential concentration issues [53].